OpenStudy (twizttiez):

What is the exact volume of a cylinder that has a height of 1.8 ft and a radius of 6.1 ft? ft3

OpenStudy (hockeychick23):

volume= πr^2(h) volume=π (6.1^2)(1.8)= 210.42

OpenStudy (twizttiez):

Can you continue to help me?

OpenStudy (hockeychick23):

sure

OpenStudy (twizttiez):

Thank you